The order reads: "Over the next 4 hours, infuse 500 ml of 5% Dextrose in Normal Saline. Add 20 MEq of KCl to solution." You know that the IV tubing set is calibrated to deliver 10gtt/mL. In drops per minute, what is the rate of flow?
Answer #1

Ans) Given: Volume to be infused- 500ml of 5% dextrose in normal saline.

20MEq of KCL added in IV 5% Dextrose

Time in which the amount to be infused- 4hrs= 4×60=240mins

Drop factor- 10gtts/ml

Formula: Flow rate= Volume (ml)/ time in mins × drop factor

Let x be the rate of flow

x= 500/240 × 10

x= 20.8 = 21ml/hr

The rate of flow is 21ml/hr.